OpenStudy (anonymous):

If f(x) = 1/3x − 2 and g(x) = −2x − 1, find (f • g)(9) 19 −18 0 −19

OpenStudy (anonymous):

what's g(9)?

OpenStudy (anonymous):

thats what I'm supposed to find I think

OpenStudy (anonymous):

its the function

OpenStudy (anonymous):

no you are supposed to find f(g(9)) ~ so we start with g(9) substitute x = 9 for function g(x) first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got -19

OpenStudy (anonymous):

tell me what you get :D *next step* after you get g(9), put the new value for x in f(x)

OpenStudy (anonymous):

-19

OpenStudy (anonymous):

yup now x = -19 for f(x)

OpenStudy (anonymous):

okay so I got the right answer

OpenStudy (anonymous):

can you help me with some more ?

OpenStudy (anonymous):

ok o.o

OpenStudy (anonymous):

If f(x) = 3x − 2 and g(x) = 2x − 1, find (f − g)(2). 7 1 −7 −1

OpenStudy (anonymous):

(f - g) (x) = 3x - 2 - (2x - 1) = 3x - 2 - 2x + 1 Substitute x = 2

OpenStudy (anonymous):

I got 1 is that correct

OpenStudy (anonymous):

3x - 2 - 2x + 1 = x - 1 = 2 - 1 = 1 yea. by the way i think i gave you the wrong explaination for the first one ._. but if you get -19, you are right lol

OpenStudy (anonymous):

If f(x) = 1/3x − 2 and g(x) = −2x − 1, find (f • g)(9) suppose to be (1/3 x - 2)(-2x-1) = -1/3 (x - 6)(2x + 1) -1/3 (9 - 6)(18 + 1) = - 19

OpenStudy (anonymous):

haha okay thanks

OpenStudy (anonymous):

If f(x) = 3x + 1 and g(x) = 2x − 4, find f[g(2)]. 25 4 0 1

OpenStudy (anonymous):

g(2) first then f(g(2)) :x 2(2) - 4 = ? then 3(?) + 1

OpenStudy (anonymous):

I got 0 for 2(2)-4

OpenStudy (anonymous):

yup then 3(0) + 1

OpenStudy (anonymous):

okay so the answer would be 1

OpenStudy (anonymous):

If f(x) = −3x − 3 and g(x) = 2x + 1, find f/g (−2). −3 3 −1 1

OpenStudy (anonymous):

hmm \[\frac{-3x-3}{2x+1}\]then sub x = -2

OpenStudy (anonymous):

i got 3 for -3(-2)-3

OpenStudy (anonymous):

and -3 for 2(-2)+1

OpenStudy (anonymous):

then 3/(-3)~

OpenStudy (anonymous):

so the answer would be -1 right

OpenStudy (anonymous):

If f(x) = 1/2x + 2 and g(x) = −3x − 1, find g[f(4)]. −19 11 −13 −11

OpenStudy (anonymous):

find f(4) then g(f(4))

OpenStudy (anonymous):

I got -13 is that correct

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1). 2 4 −2 −8

OpenStudy (anonymous):

2x + 1 - 2x - 3 x = -1

OpenStudy (anonymous):

I got -2 thats correct

OpenStudy (anonymous):

yup, gotta brb

OpenStudy (anonymous):

ok thanks for help though

OpenStudy (anonymous):

I appreciate it

OpenStudy (anonymous):

i demand medals D: